# Bones of Contention: Exploring Query-Efficient Attacks Against Skeleton Recognition Systems

### v1

@@ IV Methodology
In this section, we introduce ISAAC-K and ISAAC-N in detail. ISAAC-K can be boiled down to key joint extraction and perturbation optimization. The optimization strategy is customized to preserve naturalness of the adversarial skeletons. Moreover, ISAAC-N explores the possibility of query-free no-box attacks by replacing part-wise non-semantic joints.
@@ IV-A Key Joint Extraction
Key joints refer to those closely related to higher-level features that can directly affect the classification results. To reduce the perturbations added to the motion and improve efficiency, we need to extract the key joints for the victim model. Since the attacker has no access to the inner features of the model, we circuitously harness a local white-box model to extract key joints and transfer them to the black-box model based on the fact that the deep features learned from different models have certain transferability [35]. Here we draw on Grad-CAM [29] and adapt it to skeleton data. We obtain the significance score corresponding to the weights of different classes on the feature map based on the gradient information of the feature map before the model output layer. We calculate the weight of the $k$-th feature map corresponding to class $c$ at the $l$-th layer:
$\alpha_{k}^{l,c}=\frac{1}{N}\sum_{i=1}^{N}\frac{\partial y^{c}}{\partial F_{k,i}^{l}},$ (2)
@@
where $y^{c}$ denotes the confidence score of class $c$, $F_{k,i}^{l}$ represents the feature at the $i$-th joint in the $k$-th feature map at the $l$-th layer. Then, the significance score is calculated as:
@@
${L^{c}}[l,i]=\mathrm{ReLU}\left({\sum\limits_{k}{\alpha_{k}^{l,c}F_{k,i}^{l}\left({\mathbf{x},A}\right)}}\right),$ (3)
@@
where $\mathrm{ReLU}$ represents the ReLU activation function, and $A$ is the adjacency matrix. We rank the significance score of joints from high to low, and select the first $N_{k}$ joints as the key joints. We hope that sparse perturbations are applied only to the same joints in each frame, since it would be more practical in real-world scenarios. Though Grad-CAM may output different key joints among different frames, we use the top $N_{k}$ joints sorted in descending order of count for all key joints extracted from $T$ frames. We conduct a pilot experiment to verify the minor frame-wise key-joint difference. Given 100 videos from NTU60 RGB+D [36] and a pre-trained ST-GCN [37], we calculate the average cosine similarity of the one-hot version of key joints (e.g., $\left[0,1,1,0,0,1,...,1\right]^{T}$) from different frames. The result turns out to be 0.9719, indicating that it is convincing to use the count-based key joint selection strategy mentioned above.
Overall, Grad-CAM helps us find several joints that are crucial to surrogate model prediction. Perturbing these joints is more likely to mislead black-box models due to model transferability.
@@
IV-B Perturbation Optimization
It is empirically found that adversarial examples generated by gradient-based attacks are found mostly in the vertices of the $\ell_{\infty}$ ball [38], which is similar to the phenomenon in high-dimensional Gaussian that the majority of the probability mass is distributed around the spherical shell [39]. Besides, with the same radius, the $\ell_{\infty}$ ball is a superset of other $\ell_{p}$ balls. Therefore, searching on the $\ell_{\infty}$ ball can find stronger adversarial examples [40, 38]. To the best of our knowledge, RayS [25] has been proposed and verified as one of the most efficient hard-label attacks under black-box setting in the image domain, which reduces the search space from $\mathbb{R}^{d}$ to $\{-1,1\}^{d}$ by searching only from the ray direction space, where $d$ denotes the input dimension ($d=T\times N\times D$ for skeleton data). With RayS, we first propose the $\ell_{\infty}$ black-box attack against skeleton action recognition models. Instead of Equation 1, we turn to solving
$\mathop{\min}\limits_{{\mathbf{d}}\in{{\left\{{-1,1}\right\}}^{d}}}{g}\left({\mathbf{d}}\right),$ (4)
where $g\left({\mathbf{d}}\right)=\mathop{\arg\min}\limits_{r}\mathbbm{1}\left\{{f\left({\mathbf{x}+r\frac{{\mathbf{d}}}{{{{\left\|{\mathbf{d}}\right\|}_{2}}}}}\right)\neq y}\right\}$, denotes the decision boundary distance obtained by binary search. $r\in\mathbb{R}^{+}$ is a multiplier that controls the perturbation size.
Bone Length (Intra) Constraint.  It is low-efficient to restrict the joint position by the distance function in Equation 1 since additional velocity and acceleration restriction may also be necessitated to avoid artifacts. Instead, we restrict the bone length between the original skeleton and the adversarial skeleton to achieve a more effective and efficient attack. A prior work [18] has shown that perturbing bone length is more realizable for people. To concretely demonstrate the relation between parent joint and child joints in the bone length constraint, Figure 3 shows the 25 joints in a skeleton from the NTU60 RGB+D dataset [36]. We abstract all joints of the skeleton into a tree, with the root node being the central joint of the skeleton (joint 21 in the NTU dataset). Then we search recursively from the root node each time. If the change in a bone relative to its original length exceeds a clipping threshold, we modify its child joints according to the tree in Figure 3. For any adversarial joint ${{\mathbf{x}}_{i}^{{}^{\prime}}}$, its position is adjusted as follows.
$\begin{array}[]{l}{\mathbf{x}}_{i}^{{}^{\prime}}\leftarrow{\mathbf{x}}_{p}^{{}^{\prime}}+\frac{{{{\mathbf{d}}_{adv}}}}{{{{\left\|{{{\mathbf{d}}_{adv}}}\right\|}_{2}}}}{\mathbf{\tilde{d}}},\\ {\mathbf{\tilde{d}}}={\rm{Clip}}\left({{{\left\|{{{\mathbf{d}}_{adv}}}\right\|}_{2}},\left({1-{\varepsilon_{b}}}\right){{\left\|{{{\mathbf{d}}_{ori}}}\right\|}_{2}},\left({1+{\varepsilon_{b}}}\right){{\left\|{{{\mathbf{d}}_{ori}}}\right\|}_{2}}}\right),\end{array}$ (5)
where $\mathbf{x}_{p}^{{}^{\prime}}$ denotes the root joint in the adversarial skeleton, ${{\mathbf{d}}_{ori}}={{\mathbf{x}}_{i}}-{{\mathbf{x}}_{p}}$, ${{\mathbf{d}}_{adv}}={\mathbf{x}}_{i}^{{}^{\prime}}-{\mathbf{x}}_{p}^{{}^{\prime}}$. The subscript $i$ represents the $i$-th joint. $\rm{Clip}$ denotes the clip operation to limit the input, ${{\varepsilon_{b}}}$ represents the bone length constraint threshold. In short, we restrict the bone length to a reasonable range by pulling its child joints based on the linkage mechanism between child joints and root joint.
@@ IV-B Perturbation Optimization
Temporal Consistency (Inter) Constraint. We pose the temporal consistency constraint based on two reasons. 1) As skeleton data are expressed only by joint positions, minor perturbations may cause apparent flickers or anomalous variations between consecutive frames, as shown in Figure 4. The abrupt changes between adjacent frames make adversarial skeletons easily detectable. 2) The hierarchical version of RayS [25], which can improve the attack efficiency over the naive version, first splits the search direction into $2^{s}$ ($s$ denotes the stage in [25]) blocks and then changes the sign of each block simultaneously. It has no impact on images, but significantly affects the temporal consistency since the search directions of some joints in consecutive frames are totally opposite when (very frequently) the block consists of the last few joint positions in the first frame and the first few joint positions in the next frame. Therefore, not only can restricting the bone length of consecutive frames of the adversarial motion maintains temporal naturalness, but also be an improvement of RayS when applied to time series data. To be specific, we define $DTC$ to measure the degree of temporal consistency in the direction of perturbation between different frames of an adversarial motion, which is computed as follows:
$DTC=\frac{1}{T-1}\sum_{i=1}^{T-1}cosine\_similarity\left({{{\mathbf{\hat{d}}}}_{i}},{{{\mathbf{\hat{d}}}}_{i+1}}\right),$ (6)
where $\mathbf{\hat{d}}_{i}={\mathbf{x}}_{i}^{{}^{\prime}}-{{\mathbf{x}}_{i}}$ is the perturbation direction of frame $i$. We select the adversarial motion with the smallest DTC from all the adversarial motions during perturbation optimization, in order to make the perturbations in the adversarial motion as continuous as possible temporally. Note that the perturbations are added to all frames of the input motion.
@@
$g_{res}\left({\mathbf{d}}\right)=\mathop{\arg\min}\limits_{r}\mathbbm{1}\left\{{f\left({\Theta\left({\mathbf{x}+r\frac{{\mathbf{d}}}{{{{\left\|{\mathbf{d}}\right\|}_{2}}}},\mathbf{x},{\varepsilon_{b}}}\right)}\right)\neq y}\right\},$ (7)
@@
$\mathop{\min}\limits_{{\mathbf{d}}\in{{\left\{{-1,1}\right\}}^{\rho d}}}{g_{res}}\left({M\odot\mathbf{d}}\right),$ (8)

### v2

@@ IV Methodology
In this section, we introduce ISAAC-K and ISAAC-N in detail. ISAAC-K can be boiled down to key joint extraction and perturbation optimization. The optimization strategy is customized to preserve the naturalness of adversarial skeletons. Moreover, ISAAC-N explores the possibility of query-free no-box attacks by replacing part-wise non-semantic joints.
@@ IV-A Key Joint Extraction
Key joints refer to those closely related to higher-level features that can directly affect the classification results. To reduce the perturbations added to the motion and improve efficiency, we need to extract the key joints for the victim model. Since the attacker has no access to the inner features of the model, we circuitously harness a local surrogate model to extract key joints and transfer them to the black-box victim model based on the fact that the deep features learned from different models have certain transferability [jiang2019black]. Here we draw on Grad-CAM [selvaraju2017gradcam] and adapt it to skeleton data. The significance score for each class is derived from the gradients of the feature maps before the output layer. Specifically, the weight of the $k$-th feature map corresponding to class $c$ at the $l$-th layer is computed as:
$\small\alpha_{k}^{l,c}=\frac{1}{N}\sum_{i=1}^{N}\frac{\partial y^{c}}{\partial F_{k,i}^{l}},$ (2)
@@
where $y^{c}$ denotes the confidence score of class $c$, and $F_{k,i}^{l}$ represents the feature at the $i$-th joint in the $k$-th feature map at the $l$-th layer. Then, the significance score is calculated as:
@@
$\small{L^{c}}[l,i]=\mathrm{ReLU}\left({\sum\limits_{k}{\alpha_{k}^{l,c}F_{k,i}^{l}\left({\mathbf{x},A}\right)}}\right),$ (3)
@@
where $\mathrm{ReLU}$ represents the ReLU activation function, and $A$ is the adjacency matrix. Finally, joints are ranked according to their significance scores, and the top $N_{k}$ joints are selected as the key joints. We hope that sparse perturbations are applied only to the same joints in each frame, since it would be more practical in real-world scenarios. Though Grad-CAM may output different key joints among different frames, we use the top $N_{k}$ joints sorted in descending order of count for all key joints extracted from $T$ frames. We conduct a pilot experiment to verify the minor frame-wise key-joint difference. Given 100 videos from NTU60 RGB+D [shahroudy2016ntu60] and a pre-trained ST-GCN [yan2018stgcn], we calculate the average cosine similarity of the one-hot version of key joints (e.g., $\left[0,1,1,0,0,1,...,1\right]^{T}$) from different frames. The result turns out to be 0.9719, indicating that it is convincing to use the count-based key joint selection strategy mentioned above.
Overall, Grad-CAM helps us find several joints that are crucial to surrogate model prediction. Perturbing these joints is more likely to mislead black-box models due to model transferability. Note that the surrogate model serves only as a tool for extracting key joints rather than as a local white-box model used to generate transferable examples, as in traditional transfer-based attacks [papernot2016transferability]. Further analyses in Section V-H show that our method is insensitive to the choice of surrogate model when selecting key joints. Even when attacking a proprietary commercial system, using an open-source skeleton model to extract key joints is completely applicable.
IV-B Perturbation Optimization
Existing hard-label black-box attacks typically perform gradient estimation in a continuous search space using finite difference or zeroth order optimization [cheng2019query, cheng2020signopt]. However, each gradient update requires multiple rounds of calculating the decision boundary radius which is estimated through binary search. This results in a prohibitively large number of queries, making continuous-space optimization inefficient and impractical under hard-label constraints. It is empirically found that adversarial examples generated by gradient-based attacks are found mostly in the vertices of the $\ell_{\infty}$ ball [moon2019parsimonious], which is similar to the phenomenon in high-dimensional Gaussian that the majority of the probability mass is distributed around the spherical shell [bishop2006pattern]. Besides, with the same radius, the $\ell_{\infty}$ ball is a superset of other $\ell_{p}$ balls. Therefore, searching on the $\ell_{\infty}$ ball can find stronger adversarial examples [chen2020frank, moon2019parsimonious]. To the best of our knowledge, RayS [chen2020rays] has been proposed and verified as one of the most efficient hard-label attacks under black-box setting in the image domain, which reduces the search space from $\mathbb{R}^{d}$ to $\{-1,1\}^{d}$ by searching only from the ray direction space, where $d$ denotes the input dimension ($d=T\times N\times D$ for skeleton data). With RayS, we propose the $\ell_{\infty}$-based black-box attack against skeleton action recognition models. Instead of Equation 1, we turn to solving:
$\small\mathop{\min}\limits_{{\mathbf{d}}\in{{\left\{{-1,1}\right\}}^{d}}}{g}\left({\mathbf{d}}\right),$ (4)
where $g\left({\mathbf{d}}\right)=\mathop{\arg\min}\limits_{r}\mathbbm{1}\left\{{f\left({\mathbf{x}+r\frac{{\mathbf{d}}}{{{{\left\|{\mathbf{d}}\right\|}_{2}}}}}\right)\neq y}\right\}$ denotes the decision boundary distance obtained by binary search, a classical strategy employed in decision-based attacks [cheng2019query, chen2020rays]. $r\in\mathbb{R}^{+}$ is a multiplier that controls the perturbation size.
@@ IV-B Perturbation Optimization
Bone Length (Intra) Constraint.  It is low-efficient to restrict the joint position by the distance function in Equation 1 since additional velocity and acceleration restriction may also be necessitated to avoid artifacts. Instead, we restrict the bone length between the original skeleton and the adversarial skeleton to achieve a more effective and efficient attack. A prior work [tanaka2022adversarial] has shown that perturbing bone length is more realizable. To concretely demonstrate the relation between child joints and parent joints in the bone length constraint, Fig. 3 shows the 25 joints in a skeleton from the NTU60 RGB+D dataset [shahroudy2016ntu60]. We abstract all joints of the skeleton into a tree, with the root joint being the central joint of the skeleton (joint 21 in the NTU dataset). Then we search recursively from the root joint each time. If the change in a bone relative to its original length exceeds a clipping threshold, we modify its child joints according to the tree in Fig. 3. For the $i$-th adversarial joint ${\mathbf{x}}_{i}^{{}^{\prime}}$}, its position is adjusted as follows.
$\small\begin{array}[]{l}{\mathbf{x}}_{i}^{{}^{\prime}}\leftarrow{\mathbf{x}}_{p}^{{}^{\prime}}+\frac{{{{\mathbf{d}}_{adv}}}}{{{{\left\|{{{\mathbf{d}}_{adv}}}\right\|}_{2}}}}{\mathbf{\tilde{d}}},\\ {\mathbf{\tilde{d}}}={\rm{Clip}}\left({{{\left\|{{{\mathbf{d}}_{adv}}}\right\|}_{2}},\left({1-{\varepsilon_{b}}}\right){{\left\|{{{\mathbf{d}}_{ori}}}\right\|}_{2}},\left({1+{\varepsilon_{b}}}\right){{\left\|{{{\mathbf{d}}_{ori}}}\right\|}_{2}}}\right),\end{array}$ (5)
where $\mathbf{x}_{p}^{{}^{\prime}}$ denotes the parent joint corresponding to ${\mathbf{x}}_{i}^{{}^{\prime}}$}, ${{\mathbf{d}}_{ori}}={{\mathbf{x}}_{i}}-{{\mathbf{x}}_{p}}$, ${{\mathbf{d}}_{adv}}={\mathbf{x}}_{i}^{{}^{\prime}}-{\mathbf{x}}_{p}^{{}^{\prime}}$, $\rm{Clip}$ denotes the clip operation to limit the input, $\varepsilon_{b}$}} represents the bone length constraint threshold. In short, we restrict the bone length to a reasonable range by pulling its child joints based on the linkage mechanism between child joints and parent joints.
@@
Temporal Consistency (Inter) Constraint. We pose the temporal consistency constraint based on two reasons. 1) As skeleton data are expressed only by joint positions, minor perturbations may cause apparent flickers or anomalous variations between consecutive frames, as shown in Fig. 4. The abrupt changes between adjacent frames make adversarial skeletons easily detectable. 2) The hierarchical version of RayS [chen2020rays], which can improve the attack efficiency over the naive version, first splits the search direction into $2^{s}$ ($s$ denotes the stage in [chen2020rays]) blocks and then changes the sign of each block simultaneously. It has no impact on images, but significantly affects the temporal consistency since the search directions of some joints in consecutive frames are totally opposite when (very frequently) the block consists of the last few joint positions in the first frame and the first few joint positions in the next frame. Therefore, not only can restricting the bone length of consecutive frames of the adversarial motion maintain temporal naturalness, but also be an improvement of RayS when applied to time series data. To be specific, we define the Temporal Directional Consistency (TDC) to measure the frame-to-frame consistency of perturbation directions of the skeleton. The TDC (range: $[0,1]$) between two adjacent frames k1k-1 and $k$ is computed as:
$\text{TDC}^{[k]}=\frac{1}{2}*{cosine\_{similarity}}\left(\hat{\mathbf{d}}^{[k-1]},\ \hat{\mathbf{d}}^{[k]}\right)+\frac{1}{2},$ (6)
where $\hat{\mathbf{d}}^{[k]}=\mathbf{x}^{\prime[k]}-\mathbf{x}^{[k]}$ denotes the perturbation direction at frame $k$. Then we define a soft interpolation weight using an exponential decay function:
$\lambda^{[k]}=\exp(-\alpha\cdot\text{TDC}^{[k]}),$ (7)
where $\alpha$ is a large integer that ensures that $\lambda^{[k]}$ is close to 0 when $\text{TDC}^{[k]}$ reaches 1. In other words, when the two directions are already similar (i.e., $\text{TDC}^{[k]}\to 1$), the weight $\lambda^{[k]}$ becomes negligible. This property allows the update to largely retain the current perturbation. In contrast, when the directions differ significantly (i.e., $\text{TDC}^{[k]}\to 0$), the weight $\lambda^{[k]}$ is close to 1, indicating that the correction prioritizes the previous frame to avoid abrupt temporal changes. Therefore, the adversarial skeleton is updated by interpolating between the current and previous perturbation directions:
$\mathbf{x}^{\prime[k]}\leftarrow\mathbf{x}^{[k]}+\lambda^{[k]}\cdot\hat{\mathbf{d}}^{[k-1]}+(1-\lambda^{[k]})\cdot\hat{\mathbf{d}}^{[k]}.$ (8)
This update adaptively steers the current perturbation direction toward the previous one based on their alignment and reduces unnatural frame-to-frame variations in the adversarial motion.
@@
$\small g_{res}\left({\mathbf{d}}\right)=\mathop{\arg\min}\limits_{r}\mathbbm{1}\left\{{f\left({\Theta\left({\mathbf{x}+r\frac{{\mathbf{d}}}{{{{\left\|{\mathbf{d}}\right\|}_{2}}}},\mathbf{x},{\varepsilon_{b}}}\right)}\right)\neq y}\right\},$ (9)
@@
$\small\mathop{\min}\limits_{{\mathbf{d}}\in{{\left\{{-1,1}\right\}}^{\rho d}}}{g_{res}}\left({M\odot\mathbf{d}}\right),$ (10)
where $M$ denotes the flattened key joint mask matrix, $\odot$ stands for the Hadamard product, $\rho=N_{k}/N$ represents the joint sparsity. The sparse perturbation can reduce the search space to $\{-1,1\}^{\rho d}$, where the adversarial motion is selected from the vertices of parts of the $\ell_{\infty}$ ball.
IV-C ISAAC-N
As a byproduct, we find replacing part-wise non-semantic joints can also bring threats to skeletal models. Therefore, we propose ISAAC-N. For non-key joints obtained in the key joint extraction, we believe that some of these joints are relatively insensitive to the model, i.e., adding perturbations to the non-semantic parts has a relatively small impact on the recognition results. Even from the perspective of human visual perception, the perturbations of non-semantic parts will not affect human subjective judgment, e.g., as for the action of “picking up the mobile phone and taking selfies”, we may pay more attention to whether the arm and head are raised visually than the posture or action of the lower body. If perturbing the joints of the lower body to a large extent can cause misclassification of the model, we deem that the model is unsafe in this context. This is based on the basic fact that a robust model is expected to always focus on the arm joints regardless of changes in the lower body, i.e., standing, sitting, prostrating.
Taking the NTU120 RGB+D dataset [liu2019ntu120] as an example, we find that most of the non-semantic regions are in the lower body, which is also consistent with human cognition since most of the classes in this dataset are upper-body actions. We simply design four typical lower body postures, sit, crouch, kneel, and half kneel, to replace the lower body of the input motion, followed by root joint translation and normalization. If the replaced motion is not adversarial yet, we follow the improved RayS-based perturbation optimization mentioned in Section IV-B. Note that the temporal consistency constraint helps to improve temporal stability, which is particularly evident in ISAAC-N when large perturbations are considered. Therefore, we provide three attacking strategies for ISAAC-N: only replacing the lower body (ISAAC-NR), replacement followed by adding perturbations to only the lower body (ISAAC-NRL) and all joints (ISAAC-NRA).
V Experimental Evaluation
In this section, we begin with an experimental setup, followed by a comprehensive evaluation including quantitative and qualitative analyses of ISAAC-K and ISAAC-N, ablation study and countermeasures to show the overall performance of our proposed methods. Finally, we discuss the extension to targeted attacks and analyze the role of the surrogate model.
V-A Experimental Setup
Datasets. We use three prevalently used datasets, NTU60 RGB+D (NTU60) [shahroudy2016ntu60], NTU120 RGB+D (NTU120) [liu2019ntu120] and NW-UCLA [wang2014nw_ucla]. NTU60 was captured by three Kinect V2 cameras in 2016 and comprised of 56,880 video clips with 60 classes performed by 40 subjects. The dataset contains 40 everyday actions, 11 two-person interactions, and 9 health actions. Extended by NTU60, NTU120 was released 3 years later and was captured by the same cameras, with another 57,600 video clips and 60 classes, among which 42 classes are everyday actions, 15 classes are two-person interactions, and 3 classes are health actions. The videos were performed by 106 different subjects. NW-UCLA includes 1,494 video clips with 10 action classes. These videos were captured using three Kinect V1 cameras (20 body joints) positioned at various viewpoints. Of these views, two are used for training and one for testing. As NTU60 and NTU120 both contain 2D (17 body joints) and 3D (25 body joints) data, we use both of them in our experiments. Other datasets such as MHAD [ofli2013mhad], PKU-MMD [liu2017pkummd] are smaller Kinect-based datasets with similar characteristics. Therefore, we do not include them in our experiments. The HDM05 dataset [muller2007hdm05], released in 2007, is outdated, limited in scale, and prone to motion-capture artifacts, making it unsuitable for modern evaluations. Therefore, our experiments already validate the method’s performance under the standard and most representative data setting.
Models. We carefully select victim models to show our attack’s effectiveness and scalability. For NTU60 and NTU120, we first select six GCN (Graph Convolutional Network)-based models (MS-AAGCN [shi2020msaagcn], CTR-GCN [chen2021ctrgcn], DG-STGCN [duan2022dgstgcn], MS-G3D [liu2020msg3d], ST-GCN [yan2018stgcn] and ST-GCN++ [duan2022stgcn++]) due to their representativeness. We also involve two SOTA GCN-based models (TD-GCN [liu2023tdgcn] and LA-GCN [xu2023lagcn]) for NW-UCLA. In recent years, the emergence of Vision Transformer (ViT) has significantly enhanced classification performance in skeleton action recognition tasks. Therefore, we choose two SOTA accessible ViT-based 3D models (Hyperformer [zhou2022hypergraph] and SkateFormer [do2024skateformer]) for all datasets. We train these models from scratch unless the pre-trained models are provided publicly from a public toolbox, PYSKL [pyskl].
Competitors. We compare our method with BASAR [diao2021basar] since it is the only existing open-source attack under the black-box setting. BASAR uses a random search to find the adversarial motion in the human manifold. For a fair comparison, we modify BASAR from $\ell_{2}$ to $\ell_{\infty}$. We will provide the justification when introducing the metrics later. We do not compare our attack with the later proposed black-box attack, QESAR [kang2023qesar], for the following reasons. 1) QESAR is not open-source, and its reproduction is challenging due to insufficiently clarified constraints in its paper. 2) QESAR employs a decision-based strategy similar to BASAR, which is time-consuming due to its reliance on random search. As a result, although QESAR claims to reduce the query number compared to BASAR, it still requires an order of 1,000 queries to achieve success, whereas our attack only requires an order of 10 queries (see Section V-B).
Attack Preliminaries. For each datasets, we randomly select 500 skeletal motions. We use ST-GCN as our local surrogate model to extract key joints for NTU 60/120. The key joint number is set as 12/18 for 2D/3D data due to better overall performance. For NW-UCLA, we adopt SkateFormer as the surrogate model and select 14 key joints accordingly. We set the restriction threshold $\varepsilon$ (the $\ell_{\infty}$ norm between original and adversarial motion) as 0.4 by comprehensively considering the stealthiness-efficiency trade-off. See Section V-D for more details about different key joint numbers and different $\varepsilon$. Following [tanaka2022adversarial], we set the bone length constraint threshold $\varepsilon_{b}$ as 0.1. The decay constant $\alpha$ is set as 10 and the query limit is set as 2,000.
Metrics. We use Attack Success Rate (ASR) and Average Query (AQ) to evaluate the attack performance. ASR denotes the ratio of adversarial motions that successfully mislead the classifier. Any attack exceeding the query limit (2,000) will also be considered a failure. AQ stands for the required average query number for successful attacks. When ASR is close, lower AQ becomes more significant. To gauge the distance of joints towards root joint, we also introduce the averaged off-centered deviation $l_{c}$, which can be expressed as:
$\small{l_{c}}=\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1,j\neq c}^{N}{\left|{{{\left\|{{{\mathbf{x}}_{\left({i,j}\right)}}-{{\mathbf{x}}_{\left({i,c}\right)}}}\right\|}_{2}}-{{\left\|{{\mathbf{x}}_{\left({i,j}\right)}^{{}^{\prime}}-{\mathbf{x}}_{\left({i,c}\right)}^{{}^{\prime}}}\right\|}_{2}}}\right|}},$ (11)
where ${\mathbf{x}}_{\left({i,j}\right)}$}} denotes the $j$-th joint of the $i$-th motion, $c$ is the root joint of the skeleton we select (e.g., joint 21 in the NTU dataset) and $n$ is the number of adversarial motions. As a complement to the metrics, following previous works [zheng2020towards, diao2021basar], we use the averaged joint acceleration deviation $\Delta a$ to quantitatively measure the difference between original and adversarial motions:
$\small\Delta a=\frac{1}{nTN}\sum_{i=1}^{n}\left\|\ddot{\mathbf{x}}_{(i)}-\ddot{\mathbf{x}}^{{}^{\prime}}_{(i)}\right\|_{2},$ (12)
where $\ddot{\mathbf{x}}_{(i)}$ denotes the joint acceleration deviation of the $i$-th motion. $l_{c}$ and $\Delta a$ describe the static and dynamic changes of bones, respectively. We replace the averaged joint position deviation $l$ used in BASAR with $l_{c}$ for two reasons. On the one hand, the skeletons may be translated as a whole in the image, resulting in an inaccurate calculation of the absolute distance in $l$. On the other hand, it is unfair to compare the $\ell_{2}$ distance in $l$ since we consider $\ell_{\infty}$ restriction in our attack. Therefore, we calculate the average distance between each joint to the root joint to ensure a fair comparison as much as possible.
TABLE II: Attack performance comparison on NTU60.
Model
(Test Accuracy)
Attack NTU60(2D) NTU60(3D)